How many grams are there in 13.5 moles of NaCl? (show your work)
1
Expert's answer
2017-02-24T07:24:18-0500
Main formula: m=n*M,

n- number of moles;

M-molar mass of compound;

n(NaCl)=13.5 mol;

M(NaCl)=M(Na)+M(Cl)=23.0+35.5=58.5 (g/mol)

M(NaCl)=n(NaCl)*M(NaCl)=13.5mol* 58.5g/mol=790g

Answer: 790g